the scale on a road map is 1cm= 500 miles. if the actual distance is 1,100 miles, what is the distance on the map between the two cities
1
Expert's answer
2013-06-03T10:57:45-0400
We denote by xthe distance between two cities. if 1 cm=500 miles, then we can make the proportion: 1/500=x/1100. Whence x=2,2cm.